The Ksp of magnesium hydroxide is 1.2 x 10^-11. Find the concentration of each ion at saturation point
Answer #1

Mg(OH)2 <----> Mg2+ + 2OH-

Kps= s* (2s)2

s3=Kps/4

s= cubic root(Ksp/4)= 1.44*10-4 M

[Mg2+]= 1.44*10-4 M

[OH-]= 2.88*10-4 M